OpenStudy (anonymous):

The dwarf sea horse swims at a rate of 52.68 feet per hour.Convert this speed to inches per minut. I need help

OpenStudy (tkhunny):

Multiply by 1 as many times as necessary. \(\dfrac{52.68\;ft}{1\;hr}\cdot\dfrac{12\;in}{1\;ft}\cdot\dfrac{1\;hr}{60\;min}\)